# Local Surface Parameterizations via Geodesic Splines

### v1

@@ 5. Limitations, Future Work and Conclusion
Figure 14. In the presence of particularly large surface features, holonomy smoothing produces large rotations and shears the integration front, causing curves to be near-parallel and overlap.
The primary technical limitation of our method is that our maps are restricted to discs in tangent space, which also restricts their maximum radii. Unlike “log map-first” methods on meshes (Sharp et al., 2019; Herholz and Alexa, 2019), which assign tangent coordinates to each vertex and indirectly create a star around the origin in tangent space, our method restricts the geometry of the domain and thus cannot expand into large maps across complex surfaces without intersecting or rapidly separating (adjacent) geodesics, which lead to a poorly interpolated spline surface. Even holonomy smoothing cannot entirely ameliorate this issue, particularly around large protrusions on the surface, because it will try to emulate a surface of constant Gaussian curvature that will eventually cause adjacent geodesics to run nearly parallel with each other (Fig. 14). The parallel transport assumption along the current integration front $\psi_{j}$ used in holonomy smoothing can also degrade in accuracy as the geodesics separate over several tracing iterations, though this can be remedied by tracing more geodesics to reduce the inter-geodesic separation distances. Thus, a method that can separate varying angular density between radial geodesics, varying termination steps for each geodesic, and an interpolation method that can accurately stitch such geodesics together, is required for a fully robust “exp map-first” solution to the local parameterization problem, though our much simpler approach already produces excellent results on many complex surfaces and previously unsupported geometric representations.
@@ Appendix A Instability of Strip Holonomy
where $\mathsf{L}$ is the 1D (positive-semidefinite) Laplacian, $\mathsf{I}$ is the identity matrix, and $\bm{\Phi}_{j}$ is defined coordinate-wise by $(\bm{\Phi}_{j})_{i}=(\phi_{i-1,j}-\phi_{i,j})$. However, we typically want $\kappa$ to be small for smoothing to be effective, so we will drop the $\frac{1}{\kappa}\mathsf{I}$ term. Since $\bm{\Phi}_{0}=\bm{\Theta}_{0}=\mathbf{0}$, we have $\mathsf{L}\bm{\Theta}_{1}=\bm{\Phi}_{1}$, and then $\mathsf{L}\bm{\Theta}_{2}=\bm{\Phi}_{2}-2\bm{\Phi}_{1}$, $\mathsf{L}\bm{\Theta}_{3}=\bm{\Phi}_{3}-(2\bm{\Phi}_{2}-2\bm{\Phi}_{1})$, and in general

### v2

@@ 5. Limitations, Future Work and Conclusion
The primary technical limitation of our method is that our maps are restricted to discs in tangent space, which also restricts their maximum radii. Unlike “log map-based” methods on meshes (Sharp et al., 2019; Herholz and Alexa, 2019), which assign tangent coordinates to each vertex and indirectly create a star around the origin in tangent space, our method restricts the geometry of the domain and thus cannot expand into large maps across complex surfaces without intersecting or rapidly separating (adjacent) geodesics, which lead to a poorly interpolated spline surface. Even holonomy smoothing cannot entirely ameliorate this issue, particularly around large protrusions on the surface, because it will try to emulate a surface of constant Gaussian curvature that will eventually cause adjacent geodesics to run nearly parallel with each other (Fig. 9). However, even prior work such as the vector heat method struggles in these challenging scenarios; a challenging example from Fig. 11 is shown in Fig. 16 where both our method and the vector heat method produce distorted results. The parallel transport assumption along the current integration front $\psi_{j}$ used in holonomy smoothing can also degrade in accuracy as the geodesics separate over several tracing iterations, though this can be remedied by tracing more geodesics to reduce the inter-geodesic separation distances. Thus, a method that can separate varying angular density between radial geodesics, varying termination steps for each geodesic, and an interpolation method that can accurately stitch such geodesics together, is required for a fully robust “exp map-based” solution to the local parameterization problem, though our much simpler approach already produces excellent results on many complex surfaces and previously unsupported geometric representations.
Figure 16. A challenging example from the dataset of Myles et al. (2014) where our method’s log map triangulated from surface samples exhibits high distortion (right), with shearing near the shark’s dorsal fin, and wrapping around itself near the tail (unlike Fig. 14 which used a custom renderer, the Blender renderer used here does not support map-aware compositing). However, when the vector heat method (Sharp et al., 2019) (left) is applied to the same set of sample points, it also struggles to reproduce the dorsal fin, though it produces an injective map. The two distortion metrics from Fig. 11 are given as well, where “SD” represents symmetric Dirichlet energy, and “LSCM” represents LSCM energy.
@@
Acknowledgements.
This project was funded in part by an NSERC Discovery Grant (RGPIN-2023-05120) and an Ontario Early Researchers Award. The first author was funded by an NSERC Canada Graduate Scholarship — Doctoral. We thank Victor Rong, Honglin Chen, Silvia Sellán, and Towaki Takikawa for proofreading.
@@ Appendix A Instability of Strip Holonomy
where $\mathsf{L}$ is the 1D (positive-semidefinite) Laplacian, $\mathsf{I}$ is the identity matrix, and $\bm{\Phi}_{j}$ is defined coordinate-wise by $(\bm{\Phi}_{j})_{i}=(\phi_{i-1,j}-\phi_{i,j})$. However, we typically want $\frac{1}{\kappa}$ to be small for smoothing to be effective, so we will drop the $\frac{1}{\kappa}\mathsf{I}$ term. Since $\bm{\Phi}_{0}=\bm{\Theta}_{0}=\mathbf{0}$, we have $\mathsf{L}\bm{\Theta}_{1}=\bm{\Phi}_{1}$, and then $\mathsf{L}\bm{\Theta}_{2}=\bm{\Phi}_{2}-2\bm{\Phi}_{1}$, $\mathsf{L}\bm{\Theta}_{3}=\bm{\Phi}_{3}-(2\bm{\Phi}_{2}-2\bm{\Phi}_{1})$, and in general